OpenStudy (anonymous):

Reaner recycling shreds 7/8 tons of aluminum each day. The machines can shred 1/24 tons aluminum per cycle. How many cycles will be needed to shred the aluminum?

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

\[\huge \frac{7}{8}\div \frac{1}{24}\] in other words

OpenStudy (anonymous):

\[\huge \frac{ 7 }{ 8 }\times24=?\]

OpenStudy (anonymous):